# Copyright-Aware Incentive Scheme for Generative Art Models Using Hierarchical Reinforcement Learning

### v1

@@ 3.1. Preliminaries
$c^{t}_{i}$ Copyright loss of training sample $i$ at round $t$
@@
$sim_{sem}$ Semantic similarity
$sim_{per}$ Perceptual similarity
@@ 4.2. Modeling Copyright Loss
(14) $sim_{sem}=MSE(CLIP(s_{i})-CLIP(y_{i}))$
Perceptual Similarity
@@
Then, the copyright loss of sample $i$ is defined as:
(16) $c_{i}=asim_{sem}+bsim_{per}$
@@
Workflow Now we can compute the copyright loss with the metrics defined above for each training sample with respect to each generated image. As shown in Figure. 2, the workflow for this part is as follows: initially, each training sample is passed through the CLIP model (Radford et al., 2021) to generate captions, which are fed into a latent diffusion model to generate corresponding images. Next, we calculate the semantic similarity $sim_{sem}$ using Equation 14. Following this, in line with the framework proposed by Zhang et al. (Zhang et al., 2018), we compute the perceptual similarity $sim_{per}$. Once we have obtained both $sim_{sem}$ and $sim_{per}$, we proceed to calculate the raw copyright score $c^{t}_{i}$ for each sample $i$ at iteration round $t$. Finally, the total copyright loss for data holder $k$ at round $t$ is determined by summing the individual scores across all samples, as expressed in the equation below:
(17) $c^{t}_{k}=\sum_{i}c^{t}_{i}$
@@
19:          Compute rewards: $r_{t}=\lambda R_{con}-\delta R_{cop}$, where: $R_{con}=\sum_{k}{p_{k}}\times\hat{\beta_{k}},\quad R_{cop}=\sum_{k}{p_{k}}\times c_{k}$
@@ 4.3.1. Outer Budget Allocation
The $Q(\mathcal{M})$ is the output quality of the model (See Equation. 23). The design of the reward $R_{t}$ means that at the last round $T$, the reward will be the model quality measured by the metric 23; at the middle rounds $t<T$, we will attribute a small amount of reward to encourage the budget distribution process.
@@ 4.3.2. Inner Budget Allocation
where $R_{con}=\sum_{k}{p_{k}}\times\hat{\beta_{k}}$ and $R_{cop}=\sum_{k}{p_{k}}\times c_{k}$. The reward is based on the intuition that data holder with high contribution and high copyright should be given high payoffs. We design the reward to help the system pick the data holders with the larger contribution and the lower the copyright loss. The setting can help the system spend the budget in a smart way that allocate more payoff to the data holders that make greater contribution to the training while doesn’t need to compensate a lot for the copyright loss.
@@ 4.3.3. Workflow
    Train the model with data in all data holders. Compute the contribution of each training sample $\hat{\beta}^{t}_{i}$ and the copyright loss of each training sample $c^{t}_{i}$.

### v2

@@ 3.1. Preliminaries
$c_{i}$ Comprehensive similarity of sample i
$\hat{c^{t}_{i}}$ Comprehensive copyright loss of training sample $i$ at round $t$
@@
$Dist_{sem}$ Semantic distance
$Dist_{per}$ Perceptual distance
@@ 4.2. Modeling Copyright Loss
(14) $Dist_{sem}=MSE(CLIP(s_{i})-CLIP(y_{i}))$
Perceptual Distance
@@
Then, the comprehensive similarity of sample $i$ is defined as:
(16) $c_{i}=aDist_{sem}+bDist_{per}$
@@
We normalize and process the comprehensive similarity of each sample. The copyright loss of a sample $i$ at time $t$ is defined as follows:
(17) $\hat{c_{i}}^{t}=1-\frac{c_{i}^{t}-c_{min}^{t}}{c_{max}^{t}-c_{min}^{t}}.$
Workflow Now we can compute the copyright loss with the metrics defined above for each training sample with respect to each generated image. As shown in Figure. 2, the workflow for this part is as follows: initially, each training sample is passed through the CLIP model (Radford et al., 2021) to generate captions, which are fed into a latent diffusion model to generate corresponding images. Next, we calculate the semantic distance $Dist_{sem}$ using Equation 14. Following this, in line with the framework proposed by Zhang et al. (Zhang et al., 2018), we compute the perceptual distance $Dist_{per}$. Once we have obtained both $Dist_{sem}$ and $Dist_{per}$, we proceed to calculate the comprehensive similarity $\hat{c_{i}}^{t}$ for each sample $i$ at iteration round $t$. Finally, the total copyright loss for data holder $k$ at round $t$ is determined by summing the individual scores across all samples, as defined in the equation below:
(18) $c^{t}_{k}=\sum_{i}\hat{c^{t}_{i}}$
@@
19:          Compute rewards: $r_{t}=\lambda R_{con}-\delta R_{cop}$, where: $R_{con}=\sum_{k}{p_{k}^{t}}\times\hat{\beta_{k}}^{t},\quad R_{cop}=\sum_{k}{p_{k}^{t}}\times\hat{c_{k}}^{t}$
@@ 4.3.1. Outer Budget Allocation
The $Q(\mathcal{M})$ is the output quality of the model (See Equation. 24). The design of the reward $R_{t}$ means that at the last round $T$, the reward will be the model quality measured by the metric 24; at the middle rounds $t<T$, we will attribute a small amount of reward to encourage the budget distribution process.
@@ 4.3.2. Inner Budget Allocation
where $R_{con}=\sum_{k}{p_{k}^{t}}\times\hat{\beta_{k}}^{t},\quad R_{cop}=\sum_{k}{p_{k}^{t}}\times\hat{c_{k}}^{t}$ The reward is based on the intuition that data holder with high contribution and high copyright should be given high payoffs. We design the reward to help the system pick the data holders with the larger contribution and the lower the copyright loss. The setting can help the system spend the budget in a smart way that allocate more payoff to the data holders that make greater contribution to the training while doesn’t need to compensate a lot for the copyright loss.
@@ 4.3.3. Workflow
    Train the model with data in all data holders. Compute the contribution of each training sample $\hat{\beta}^{t}_{i}$ and the copyright loss of each training sample $\hat{c_{i}}^{t}$.